OpenStudy (anonymous):

Find a formula for a quadratic function that passes through the points (2,0), (-1,0) and (0,4)

OpenStudy (anonymous):

The general quadratic function: \(y=ax^2+bx+c\). Plug in your points: \[\begin{cases}4=0a+0b+c\\0=a-b+c\\0=4a+2b+c\end{cases}\] Solve for \(a,b,c\).

OpenStudy (anonymous):

So I plug in x and y from one of the points?

OpenStudy (marissalovescats):

I believe that you know from the points that the x values are 2, -1, and 0. And you can make your binomials (remember the sings change) (we just just leave the 0 one out because it wont make a difference on the answer) (x-2)(x+1)(x+0)

OpenStudy (anonymous):

Do you just foil it out with this method?

OpenStudy (marissalovescats):

Yup!

OpenStudy (marissalovescats):

I'm pretty sure that's what they want you to do @sithsandgiggles correct me if I'm wrong

OpenStudy (anonymous):

(x^2+x-2)(x+0) x^3 + 0 + x^2 + 0 - 2x - 0

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Kainui

OpenStudy (marissalovescats):

(x^2-x-2)(x+0) -2+1 is -1

OpenStudy (anonymous):

x^3+x^2-2x

OpenStudy (marissalovescats):

x^3-x^2-2x

OpenStudy (anonymous):

I am thinking that I have to solve three linear equations to find the formula for this quadratic. I'm not sure about your way @marissalovescats

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

look at sithsandgigles first post

OpenStudy (dan815):

he's solved it for you

OpenStudy (anonymous):

so I have to solve all three equations for a, b, and c?

OpenStudy (dan815):

all quadratic eqn can be put in the form y=ax^2+bx+c we have to find the a b and c constant that satisfy your points

OpenStudy (dan815):

yes

OpenStudy (dan815):

tell me if u get stuck on something else

OpenStudy (anonymous):

thanks dan

OpenStudy (anonymous):

1) 0 = 4a + 2b + c 2) 0 = a + -b + c 3) 4 = a + b + c (0 = 4a + 2b + c) 0 = a - b + c 0 = -3a - b

OpenStudy (anonymous):

can someone verify this is right?

OpenStudy (anonymous):

@dan815 @satellite73

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

i think it is way easier than all of that

OpenStudy (anonymous):

\[ (2,0), (-1,0) \] tells you the zeros are \(2\) and \(-1\) so it must factor as \[f(x)=a(x-2)(x+1)\]

OpenStudy (anonymous):

since \(f(0)=4\) you know \[a(0-2)(0+1)=-2a=4\] making \(a=-2\)

OpenStudy (anonymous):

your quadratic is therefore \[f(x)=-2(x-2)(x+1)\]

OpenStudy (anonymous):

I thought there was an easier way, I was doing about 20 steps with solving both 3 equations and then substituting two of the equations to eliminate c and then substituting the bottom two equations

OpenStudy (anonymous):

you would have to do that if you were not told the two zeros, but since you were, it was much easier than all that mess

OpenStudy (anonymous):

how do we know it must factor like you did? @satellite73

OpenStudy (anonymous):

we are not given the equation in ax^2+bx+c form

OpenStudy (anonymous):

@marissalovescats, that method also works, but \((0,4)\) is not a root.